OpenStudy (fibonaccichick666):

Just a hint or link to similar problem please! Solve X'=Ax+F. More below

OpenStudy (fibonaccichick666):

\[A=\left[\begin{matrix}0 & 4 \\ -1 & 0\end{matrix}\right]\] \[F=\left[\begin{matrix}e^{2t} & 0\\0 &1\end{matrix}\right]\]

OpenStudy (fibonaccichick666):

x=[x(t);y(t)]

OpenStudy (anonymous):

Are you sure about the dimensions of \(F\)?

OpenStudy (fibonaccichick666):

nope you're right

OpenStudy (fibonaccichick666):

F=[sint ; 0]

OpenStudy (anonymous):

Okay, yeah I was wondering how a 2x2 would work on the RHS... Anyway, this seems like a fairly standard undetermined-coefficients problem.

OpenStudy (fibonaccichick666):

I know A has complex eigenvalues

OpenStudy (anonymous):

What are you getting for your homogeneous solution?

OpenStudy (fibonaccichick666):

\[X'= c_1e^{2it}v_1+c_2e^{-2it}v_2\]

OpenStudy (fibonaccichick666):

v1 and v2 are the repective eigenvectorss

OpenStudy (anonymous):

Okay, looks right. Euler's formula will help with the rewriting of the solution. As for the nonhomogeneous part: have you worked out any undetermined-coefficients approaches before for systems?

OpenStudy (fibonaccichick666):

Can do the rewrite, but issue is the other part. No I haven't :/

OpenStudy (fibonaccichick666):

Also upon rewrite I got \[x=c_1v_1(cos(2t)+isin(2t))+c_2v_2(cos(-2t)+isin(-2t))\]

OpenStudy (anonymous):

The idea is the same as for single ODEs. Say the nonhomogeneous function is something like \(ae^t\). Then we might guess a solution of the form \(\vec{a}e^t\). \[{\bf{x}}'-\begin{pmatrix}0&4\\-1&0\end{pmatrix}{\bf{x}}=\begin{pmatrix}1\\0\end{pmatrix}\sin t\] For a regular ODE we might consider a trial solution of \(a\sin t+b\cos t\), so here we'll try \(\vec{a}\sin t+\vec{b}\cos t\). So, if our trial solution is \({\bf x}_p\), then \[{\bf x}_p=\vec{a}\sin t+\vec{b}\cos t~~\implies~~{{\bf x}_p}'=\vec{a}\cos t-\vec{b}\sin t\] Substituting into the original system, we have \[\left(\vec{a}\cos t-\vec{b}\sin t\right)-\begin{pmatrix}0&4\\-1&0\end{pmatrix}\left(\vec{a}\sin t+\vec{b}\cos t\right)=\begin{pmatrix}1\\0\end{pmatrix}\sin t\] From here we'd simplify and solve for \(\vec{a}\) and \(\vec{b}\).

OpenStudy (fibonaccichick666):

let me try

OpenStudy (fibonaccichick666):

we can't do the matrix multiplication though?

OpenStudy (anonymous):

Why's that? We have a 2x2 matrix multiplied by a 2x1 vector. We would end up with another 2x1 vector.

OpenStudy (fibonaccichick666):

ohhh cause it's a vector

OpenStudy (fibonaccichick666):

duh

OpenStudy (fibonaccichick666):

k so I get \[[acost~~~-bsint]-[4bcost~~~-asint]=[1;0]sint\] Is that right?

OpenStudy (fibonaccichick666):

so b= 1/3?

OpenStudy (anonymous):

You're half-right... Keep in mind that \(\vec{b}=\begin{pmatrix}b_1\\b_2\end{pmatrix}\). What does it means for \(\vec{b}=\dfrac{1}{3}\)?

OpenStudy (fibonaccichick666):

b1=1 b2=3?

OpenStudy (anonymous):

No, you found one component of \(\vec{b}\), but we're still missing the other. Here's my approach: \[\begin{align*} \left(\vec{a}\cos t-\vec{b}\sin t\right)-\underbrace{\begin{pmatrix}0&4\\-1&0\end{pmatrix}}_{A}\left(\vec{a}\sin t+\vec{b}\cos t\right)&=\begin{pmatrix}1\\0\end{pmatrix}\sin t\\\\ \left(\vec{a}-A\vec{b}\right)\cos t-\left(\vec{b}+A\vec{a}\right)\sin t&=\begin{pmatrix}1\\0\end{pmatrix}\sin t \end{align*}\] which gives us two equations: \[\vec{a}-A\vec{b}=\begin{pmatrix}0\\0\end{pmatrix}\quad\text{and}\quad \vec{b}+A\vec{a}=\begin{pmatrix}1\\0\end{pmatrix}\] where \(\vec{a}\) and \(\vec{b}\) are 2x1 vectors.

OpenStudy (anonymous):

Solving for \(\vec{a}\) in the first equation, we can substitute into the second to solve for \(\vec{b}\): \[\vec{a}=A\vec{b}~~\implies~~\vec{b}+A^2\vec{b}=(I_2+A^2)\vec{b}=\begin{pmatrix}1\\0\end{pmatrix}~~\implies~~\vec{b}=(I_2+A^2)^{-1}\begin{pmatrix}1\\0\end{pmatrix}\] Computing gives \(\vec{b}=\begin{pmatrix}1/3\\0\end{pmatrix}\).

OpenStudy (anonymous):

You can then plug this back into the first equation and solve for \(\vec{a}\), and you'll have your nonhomogeneous solution.

OpenStudy (fibonaccichick666):

ok, and was my rewrite using euler's right?

OpenStudy (anonymous):

Yes, provided that \(v_1\) and \(v_2\) are correct. You can do some simplification to have a homogeneous solution in terms of \(C_1v_1\sin t+C_2v_2\cos t\).

OpenStudy (fibonaccichick666):

and isn't a just [-1/3; 0]?

OpenStudy (anonymous):

Close, you have your components switched around.

OpenStudy (fibonaccichick666):

gosh darnit

OpenStudy (anonymous):

Here's another example worked out if you need some more practice: http://tutorial.math.lamar.edu/Classes/DE/NonhomogeneousSystems.aspx

OpenStudy (fibonaccichick666):

Thank you. Yea, I need more practice to say the least

OpenStudy (anonymous):

No problem!

OpenStudy (fibonaccichick666):

I'm so screwed for this test

OpenStudy (fibonaccichick666):

this is the review sheet, and we have not really learned probably 75% of it

OpenStudy (anonymous):

Well if you have any other questions, there are at least a few of us here that can help you out. It's possible that undetermined coefficients might not have been the intended approach. There might be a different way using elimination techniques.

OpenStudy (fibonaccichick666):

there is no guide to how to approach it. you are probably right

OpenStudy (fibonaccichick666):

the v1 and v2 are \[v_1=[-2i;1] ~~and~~v_2=[2i;i]\] right?

OpenStudy (anonymous):

Yeah those are correct.

OpenStudy (fibonaccichick666):

2i,1*****

OpenStudy (fibonaccichick666):

wait, was the i right?

OpenStudy (anonymous):

The first one, the eigenvectors are \(\pm\begin{pmatrix}2i\\1\end{pmatrix}\).

OpenStudy (anonymous):

If you're interested in another approach, it's very doable. Possibly more to your liking.

OpenStudy (fibonaccichick666):

ok thats what I thought. No this is fine. He has a tendency to make the test identical to the review sheet. I will learn the crap out of this

OpenStudy (fibonaccichick666):

thanks again sithsies

OpenStudy (anonymous):

Alright then, good luck. You're welcome!

OpenStudy (fibonaccichick666):

so doesn't the solution have to be real?

OpenStudy (fibonaccichick666):

@zepdrix I still have a few questions on this one because it is complex, can you help?

OpenStudy (fibonaccichick666):

nvm, I figured it out! :D